Kyky232:

solve the following with the substitution method y^2=x^2-9 2y=x-3

Vocaloid:

It says to use the substitution method Taking the second equation 2y=x-3, dividing both sides by 2 gives us y = (1/2)(x-3) Substitute this for y into the first equation. Then solve for x. Then go back to y = (1/2)(x-3) and calculate y.